OpenStudy (anonymous):

integrate: ( 1/sqr(1-(e^(-x)))) dx, substitute: e^(-x)=u^2

OpenStudy (dumbcow):

ok after substituting you get \[-e^{-x} dx = 2u du\] \[dx = -\frac{2u du}{e^{-x}} = -\frac{2u du}{u^{2}} = -\frac{2 du}{u}\] \[\rightarrow -2 \int\limits \frac{du}{u \sqrt{1-u^{2}}}\] now let u = cos(theta) \[du = -\sin \theta d \theta\] \[\rightarrow 2 \int\limits \sec \theta d \theta = 2 \ln (\sec \theta +\tan\theta)\]

OpenStudy (anonymous):

what is u =sin(theta) .. why did you choose cos(theta)?

OpenStudy (anonymous):

*what if u is sin(theta)..?

OpenStudy (dumbcow):

@hpfan5 sorry i left without logging off by accident anyway, you are right you could sub either sin or cos for u and it would still work i chose "u = cos" because integral of sec is easier than integral of csc

OpenStudy (anonymous):

ok thank you :)